OpenStudy (anonymous):

solve each of the equations for 0≤x<2π. 2cos^2 x-1-cosx=0 sin^2 x+cos2x-cosx=0

hero (hero):

Pretty sure you can just solve this as a quadratic

OpenStudy (anonymous):

1st equation 2cos^2(x)-cos(x)-1=0 (2cos(x)+1)(cos(x)-1)=0 2nd equation sin^2(x)+cos^2(x)-cos(x)=0 1-cos^2(x)+cos^2(x)-cos(x)=0 1-cos(x)=0 -(cos(x)-1)=0 similar solution cos(x)-1=0 cos(x)=1 x=cos^-1(1) x=90 degrees or pi/2